# Extending the SAREF4ENER Ontology with Flexibility Based on FlexOffers

### v2

@@ I Introduction
The last decades have seen an increasing usage of renewable energy, and the development of new technologies and concepts to enable its exploitation. In particular, among those new concepts, one has crucial importance for handling the intermittency that comes with renewable energy generation: flexibility [1], i.e., the capability of changing energy loads in time or amount of energy. Flexibility enables optimizing energy consumption to achieve several objectives: renewable energy usage maximization [2], peak shaving [3, 4], CO2 emissions minimization [5], energy cost minimization [6, 7], provision of ancillary services [8], congestion prevention [9], and demand response [10]. Several models have been developed for flexibility. Many of them are accurate in their description of flexibility, but scale poorly with respect to the length of time horizons and the number of devices. A model that solves these issues is FlexOffers [11], a device-independent model that represents flexibility by using energy constraints that approximate the amount of available flexibility, but scales very well in scenarios with many devices or long time horizons. FlexOffers generates flexoffers (FOs) objects, which describe flexibility from the considered devices.FlexOffers is a powerful model used in more than 30 European and national projects, whose real-life implementation has been advancing in the last decade. An important aspect is its interaction with smart devices/buildings and the ontologies representing them. To date, a widely used ontology for smart buildings is the Smart Application REFerence (SAREF) [12] ontology. Specifically, several extensions of SAREF have been created for different domains: SAREF for Energy (SAREF4ENER) [13] has been designed to represent energy devices and services, and is the de facto standard for that. However, SAREF4ENER fails to capture energy flexibility for many use cases, specifically when flexibility constraints become more complex: this is the case for advanced devices such as batteries, electric vehicles (EVs), and heat pumps. This raises the need for an ontology capable of handling these cases. To meet this challenge, an ontology called domOS Common Ontology (dCO [^1] https://w3id.org/dco[14] has been proposed to ensure full flexibility support in the interoperability of smart buildings and services. The contribution we bring in this paper is the dCO flexibility module: this module extends SAREF4ENER in its representation of flexibility services, enabling compatibility with the full Flexoffers model and constraints, and ensuring that the most relevant flexible devices and use cases for them are represented. In this paper, we show how the flexibility module of dCO extends SAREF4ENER, how it represents the full Flexoffer model and constraints, and some advanced use cases that can be represented by it. The remainder of this paper is organized as follows. Section II describes FOs and the SAREF4ENER and dCO ontologies, while Section III shows how dCO extends SAREF4ENER and enables full FO representation. Section IV evaluates the dCO flexibility module, showing which use cases can and cannot be represented by FOs and SAREF4ENER. Finally, Section V concludes the work and provides ideas for future research.
@@ II-A Running Example
We start by making a running example so that we can describe the Flexoffer concept in the following subsections. Here, we will consider the case of a heat pump operating in a room. The heat pump is a Daikin Altherma [15], with maximum operational power $P_{max}=4.6$ kW and coefficient of performance (COP) $3.65$. This model supports the SG-Ready interface [16, 17], a standardized smart grid communication protocol that enables seamless integration with modern energy management systems. The room has a rectangular section with sides of $4$ and $5$ meters respectively, and it is $3$ meters high. The heat transfer coefficient is $c_{ht}=6\frac{W}{m^{2}K}$ [18], and heat is dispersed through one of the $3$m x $4$m side walls. The room temperature is $295$K ($22^{\circ}$), while the outside temperature is $275$K ($2^{\circ}$). We assume that room temperature must stay between $293$K ($20^{\circ}$) and $297$K ($24^{\circ}$), respectively, because of user settings. For reference, the amount of energy (heat) needed to maintain the temperature to $295$K for one hour can be obtained by
@@ II-B FlexOffer Life Cycle
The Flexoffers model describes flexibility in a unified format for different device types [11, 19], by providing an approximated representation of flexibility. In exchange for the small amount of accuracy lost with respect to exact models, modeling flexibility with FOs allows for much better scalability over long time horizons and high numbers of devices. FOs provide constraints for energy consumption: optimizing an FO means solving an optimization problem on energy consumption (e.g., cost minimization) within the constraints given by the FO. The solution of this optimization problem is a set of values $(e_{1},\ldots,e_{T})$ for energy consumption at each time unit: this is called a schedule. FOs can also be aggregated: this means that, from $M$ FOs, it is possible to generate $N\ll M$ FOs that represent their combined flexibility, with some losses. The opposite process is called disaggregation: from a schedule for an aggregated FOs, it is possible to generate schedules for the original FOs so that their combined energy consumption is the same as the original schedule.
The FO life cycle is shown in Figure 1. Two main actors are involved: an automatic agent representing the prosumer/device owner, which creates and executes the FOs, and another one who manages and optimizes the FOs, which is often either the Balance Responsible Party (BRP) or the aggregator: for clarity, in this paper we will assume it to be the aggregator [20]. FOs are first generated by the prosumer agent from previous forecasts and analysis, and then sent to the aggregator, who decides whether to accept the FO or not and notifies its decision to the prosumer. If the FO is rejected, the process ends here; if it is accepted, the FO is aggregated to other FOs and optimized with them. After this, a schedule is determined for the FO and sent back to the prosumer agent for execution.
Figure 1: A schematic description of the FOs life cycle (taken from [21]).
@@ II-C Standard FlexOffers
The simplest FO constraints are start time constraints and slice (energy) constraints. We are considering time as discrete and divided into time units, also called time slices: their duration is one hour in the running example. Start time constraints define the earliest and latest time at which the load can start. Energy constraints define, at each time unit, the minimum and maximum amount of energy that can be consumed from the considered energy load: in other words, at each time unit $t$, two numbers $emin_{t}$ and $emax_{t}$ are defined such that $emin_{t}\leq e_{t}\leq emax_{t}$, where $e_{t}$ is the energy consumption during time unit $t$. An FO with just these constraints is called a Standard FlexOffer (SFO). We can see an SFO example in Figure 2(a) for the running example for eight time units. The pink bars indicate the amount of energy that can be consumed at each time unit: here, at each time unit, it varies between 0.303 kWh and 0.478 kWh. At each time, the horizontal bars indicate a possible amount of energy consumption $e_{t}$.
@@ II-E Dependency FlexOffers
A dependency energy constraint defines dependencies between time slices. At each time unit $t$, a lower and an upper bound on the amount of consumable energy are defined, depending on how much energy has been consumed until time unit $t-1$. Formally, this can be expressed as $a\cdot(e_{1}+\ldots+e_{t-1})+b\cdot e_{t}\leq c$, where $a,b,c$ are real numbers. An FO with dependency energy constraints is called a Dependency FlexOffer (DFO). We can see an example in Figure 2(c), for the first four time units of the running example: here, at each time unit, the $x$ axis represents the amount of energy consumed up to that time, while the $y$ axis represents the amount of consumable energy at that time. We can see that, in general, the more energy has been consumed up to time $t$, the less energy can be consumed at time $t$, and vice versa.
@@ II-F Uncertain FlexOffers
Uncertain FOs (UFOs) [21] are a type of FOs that consider uncertainty from flexibility devices behavior. Two main types of uncertainty are considered: time and amount uncertainty. At each time unit, time uncertainty refers to the probability for the device to be operative at that time, while amount uncertainty refers to the amount of energy that can be consumed by the device at that time. A UFO is created in two steps. First, uncertainty related to the state (in the running example, temperature) is modeled at each time $t$; second, the probability for each energy value at each time to be feasible is calculated, taking into account both types of uncertainty. Some functions $\{f_{1},\ldots,f_{T}\}$ describing those probabilities will be obtained: those functions will define the UFO. UFOs can be visualized by choosing a probability threshold $P_{0}$. At each time $t$, the energy values having probability at least $P_{0}$ of being available for consumption can be described by intervals. Figure 2(d) shows this visualization for the running example: if we choose $P_{0}=1$, at each time unit, the available energy values are described by the pink bars, while if we choose $P_{0}=0.8$, the available energy values are described by the combined pink and blue bars. This is because at the first time unit, the status of the device is known, and the amount of available flexibility can be determined exactly, while for the next time units, there is more uncertainty (represented here by the blue bars) and the amount of flexibility certainly available (pink bars) is lower.
@@ II-G SAREF and SAREF4ENER
Smart Applications REFerence (SAREF) [^2] https://saref.etsi.org/ [12] is an ontology for IoT devices and solutions published by ETSI [^3] https://www.etsi.org/ in a series of Technical Specifications initially released in 2015 and updated in 2017. Even though its initial objective was to build a reference ontology for appliances relevant to energy efficiency, SAREF can also serve as a reference model to enable better integration of data from various vertical domains in the IoT. SAREF has been extended to different domains: specifically, SAREF for Energy Flexibility (SAREF4ENER) [13] is its extension for the energy domain, and aims to provide a standardized way to describe and interoperate energy-related devices, systems, and services. It allows us to describe devices such as solar panels, electric vehicles, and smart meters, as well as systems like microgrids, and energy management/demand response services. By using SAREF4ENER, developers can create applications that easily integrate and communicate with different energy devices, systems, and services, regardless of the manufacturer or protocol used, which enables more efficient and sustainable energy management.
@@
The World Wide Web Consortium (W3C) has developed a standard for enabling Internet of Things (IoT) interoperability, whose name is W3C Web of Things (WoT) [22]. In this context, Thing Descriptions (TDs) [23] have been developed: they describe the metadata and interfaces of the considered devices that participate in the WoT. Things are organized in TD Directories (WoT TDDs), which make it possible to register, manage, and search TDs. An important feature for IoT consumers is to send queries to the WoT directory: this is achieved by having a common ontology to annotate TDs; however, before dCO, no ontologies were integrating WoT into smart buildings. dCO [14] has been created with the purpose of achieving this integration: in this paper, we will focus on the aspect of dCO that strictly concerns energy flexibility. A description of the modules can be found in Appendix -A). The contribution of our work is the energy flexibility module, which is not described in [14]: in this paper, we explain how it extends SAREF4ENER.
@@
This section shows how SAREF4ENER has been extended to describe FOs. We show how we designed the dCO flexibility module and how we encode attributes and constraints. Details on FO messages can be found in Appendix -B.
@@ III-A FlexOffer Message
As mentioned in Section II, FOs are sent between actors: this is achieved by FO messages, which encode FO metadata and information about energy constraints. The attributes of an FO message describe the FO ID, state (e.g., accepted, rejected), time granularity, geographical location, time limits for acceptance and assignment, constraints on energy, time and price, and an energy schedule. Of these, the only ones to have an equivalent in SAREF4ENER are state, simple energy constraints (slice, TECs), price constraints and duration, while more advanced energy constraints (for DFOs and UFOs) and other attributes had to be added as extensions. Further information can be found in Appendix -B, in Table II for the attributes and Table I for the mapping to SAREF4ENER. In the next section, we show which classes and attributes we created for dCO, in order to model the parts missing in SAREF4ENER.

### v3

@@ I Introduction
The last decades have seen an increasing usage of renewable energy, and the development of new technologies and concepts to enable its exploitation. In particular, among those new concepts, one has crucial importance for handling the intermittency that comes with renewable energy generation: flexibility [LI2021100054], i.e., the capability of changing energy loads in time or amount of energy. Flexibility enables optimizing energy consumption to achieve several objectives: renewable energy usage maximization [scheller], peak shaving [FOTEINAKI2020109804, CHEN2021110932], CO2 emissions minimization [HARRIS20141], energy cost minimization [Mugnini2021, MAROTTA2021102392], provision of ancillary services [saele], congestion prevention [Fonteijn], and demand response [CHEN2018125]. Several models have been developed for flexibility. Many of them are accurate in their description of flexibility, but scale poorly with respect to the length of time horizons and the number of devices. A model that solves these issues is FlexOffers [DBLP:conf/smartgridcomm/PedersenSN18], a device-independent model that represents flexibility by using energy constraints that approximate the amount of available flexibility, but scales very well in scenarios with many devices or long time horizons. FlexOffers generates flexoffers (FOs) objects, which describe flexibility from the considered devices.FlexOffers is a powerful model used in more than 30 European and national projects, whose real-life implementation has been advancing in the last decade. An important aspect is its interaction with smart devices/buildings and the ontologies representing them. To date, a widely used ontology for smart buildings is the Smart Application REFerence (SAREF) [SAREF] ontology. Specifically, several extensions of SAREF have been created for different domains: SAREF for Energy (SAREF4ENER) [SAREF4ENER] has been designed to represent energy devices and services, and is the de facto standard for that. However, SAREF4ENER fails to capture energy flexibility for many use cases, specifically when flexibility constraints become more complex: this is the case for advanced devices such as batteries, electric vehicles (EVs), and heat pumps. In addition, there are only few other ontologies representing energy flexibility. One of them is EFOnt [Comparison1], a semantic ontology focused on energy flexibility at the building level and its applications. Another is the MAS2TERING ontology [Comparison2], which aims to integrate the Universal Smart Energy Framework (USEF) standards with the context of multi-agent systems to support the interoperability of smart grid management agents. Finally, the BIPV-BES-BEF ontology [Comparison3] targets photovoltaic, battery, and building operation and maintenance, and includes a class that describes flexible control strategies associated with buildings. However, none of these ontologies describe the flexibility constraints of the different types of devices and, like SAREF4ENER, they fail to properly represent flexibility in the aforementioned use cases. This raises the need for an ontology capable of handling these cases. To meet this challenge, an ontology called dCO  [^1] https://w3id.org/dco [DBLP:conf/esws/LaadharTP22] (d is the first letter in the EU project name, removed for double-blind peer review, and CO stands for Common Ontology) has been proposed to ensure full flexibility support in the interoperability of smart buildings and services. The contribution we bring in this paper is the dCO flexibility module, which extends SAREF4ENER to better represent all aspects of energy flexibility and enable more powerful flexibility services. It enables compatibility with the full Flexoffers model and constraints, ensuring that the most relevant flexible devices and use cases for them are supported. This is valuable because the resulting ontology, dCO, allows developers to create applications that offer: i) a more accurate representation of flexibility compared to the current state of the art, ii) applicability to a wider range of devices, such as batteries, EVs, and heat pumps, and iii) support for the fast load optimization and (dis)aggregation provided by FlexOffers.
In this paper, we show how the flexibility module of dCO extends SAREF4ENER, how it represents the full Flexoffer model and constraints, and some advanced use cases that can be represented by it. The remainder of this paper is organized as follows. Section II describes FOs and the SAREF4ENER and dCO ontologies, while Section III shows how dCO extends SAREF4ENER and enables full FO representation. Section IV evaluates the dCO flexibility module, showing which use cases can and cannot be represented by FOs and SAREF4ENER. Finally, Section V concludes the work and provides ideas for future research.
@@ II-A Running Example
We start by making a running example so that we can describe the Flexoffer concept in the following subsections. Here, we will consider the case of a heat pump operating in a room. The heat pump is a Daikin Altherma [DBLP:conf/eenergy/LilliuPS23], with maximum operational power $P_{max}=4.6$ kW and coefficient of performance (COP) $3.65$. This model supports the SG-Ready interface [DBLP:conf/isgteurope/FischerTS18, DBLP:conf/eenergy/BrusokasPSZC21], a standardized smart grid communication protocol that enables seamless integration with modern energy management systems. The room has a rectangular section with sides of $4$ and $5$ meters respectively, and it is $3$ meters high. The heat transfer coefficient is $c_{ht}=6\frac{W}{m^{2}K}$ [JAYAMAHA1996399], and heat is dispersed through one of the 33m x 44m side walls. The room temperature is 295295K ($22^{\circ}$C), while the outside temperature is 275275K ($2^{\circ}$C). We assume that room temperature must stay between 293293K ($20^{\circ}$C) and 297297K ($24^{\circ}$C), respectively, because of user settings. For reference, the amount of energy (heat) needed to maintain the temperature to 295295K for one hour can be obtained by
@@ II-B FlexOffer Life Cycle
The Flexoffers model describes flexibility in a unified format for different device types [DBLP:conf/smartgridcomm/PedersenSN18, Lilliu2021], by providing an approximated representation of flexibility. In exchange for the small amount of accuracy lost with respect to exact models, modeling flexibility with FOs allows for much better scalability over long time horizons and high numbers of devices. FOs provide constraints for energy consumption: optimizing an FO means solving an optimization problem on energy consumption (e.g., cost minimization) within the constraints given by the FO. The solution of this optimization problem is a set of values $(e_{1},\ldots,e_{T})$ for energy consumption at each time unit: this is called a schedule. FOs can also be aggregated: this means that, from $M$ FOs, it is possible to generate $N\ll M$ FOs that represent their combined flexibility, with some losses. The opposite process is called disaggregation: from a schedule for an aggregated FOs, it is possible to generate schedules for the original FOs so that their combined energy consumption is the same as the original schedule.
The FO life cycle is shown in Figure 1. Two main actors are involved: an automatic agent representing the prosumer/device owner, which creates and executes the FOs, and another one who manages and optimizes the FOs, which is often either the Balance Responsible Party (BRP) or the aggregator: for clarity, in this paper we will assume it to be the aggregator [DBLP:conf/eenergy/NeupaneSP17]. FOs are first generated by the prosumer agent from previous forecasts and analysis, and then sent to the aggregator, who decides whether to accept the FO or not and notifies its decision to the prosumer. If the FO is rejected, the process ends here; if it is accepted, the FO is aggregated to other FOs and optimized with them. After this, a schedule is determined for the FO and sent back to the prosumer agent for execution.
Figure 1: A schematic description of the FOs life cycle (taken from [DBLP:conf/eenergy/LilliuPSN23]).
@@ II-C Standard FlexOffers
The simplest FO constraints are start time constraints and slice (energy) constraints. We are considering time as discrete and divided into time units, also called time slices: their duration is one hour in the running example. Start time constraints define the earliest and latest time at which the load can start. Energy constraints define, at each time unit, the minimum and maximum amount of energy that can be consumed from the considered energy load: in other words, at each time unit $t$, two numbers $emin_{t}$ and $emax_{t}$ are defined such that $emin_{t}\leq e_{t}\leq emax_{t}$, where $e_{t}$ is the energy consumption during time unit $t$. An FO with just these constraints is called a Standard FlexOffer (SFO). We can now give a formal definition of energy flexibility.
Definition 1
With the notation just used, at each time unit $t$, we define amount flexibility as $emax_{t}-emin_{t}$. Also, if the load can start between time units $t_{0}$ and $t_{1}$, we define time flexibility as $t_{1}-t_{0}$.
We can see an SFO example in Figure 2(a) for the running example for eight time units. The pink bars indicate the amount of energy that can be consumed at each time unit: here, at each time unit, it varies between 0.303 kWh and 0.478 kWh. At each time, the horizontal bars indicate a possible amount of energy consumption $e_{t}$.
@@ II-E Dependency FlexOffers
A dependency energy constraint defines dependencies between time slices. At each time unit $t$, a lower and an upper bound on the amount of consumable energy are defined, depending on how much energy has been consumed until time unit t1t-1. Formally, this can be expressed as $a\cdot(e_{1}+\ldots+e_{t-1})+b\cdot e_{t}\leq c$, where $a,b,c$ are real numbers. An FO with dependency energy constraints is called a Dependency FlexOffer (DFO). We can see an example in Figure 2(c), for the first four time units of the running example: here, at each time unit, the $x$ axis represents the amount of energy consumed up to that time, while the $y$ axis represents the amount of consumable energy at that time. We can see that, in general, the more energy has been consumed up to time $t$, the less energy can be consumed at time $t$, and vice versa.
@@ II-F Uncertain FlexOffers
Uncertain FOs (UFOs) [DBLP:conf/eenergy/LilliuPSN23] are a type of FOs that consider uncertainty from flexibility devices behavior. Two main types of uncertainty are considered: time and amount uncertainty. At each time unit, time uncertainty refers to the probability for the device to be operative at that time, while amount uncertainty refers to the amount of energy that can be consumed by the device at that time. A UFO is created in two steps. First, uncertainty related to the state (in the running example, temperature) is modeled at each time $t$; second, the probability for each energy value at each time to be feasible is calculated, taking into account both types of uncertainty. Some functions $\{f_{1},\ldots,f_{T}\}$ describing those probabilities will be obtained: those functions will define the UFO. UFOs can be visualized by choosing a probability threshold $P_{0}$. At each time $t$, the energy values having probability at least $P_{0}$ of being available for consumption can be described by intervals. Figure 2(d) shows this visualization for the running example: if we choose $P_{0}=1$, at each time unit, the available energy values are described by the pink bars, while if we choose $P_{0}=0.8$, the available energy values are described by the combined pink and blue bars. This is because at the first time unit, the status of the device is known, and the amount of available flexibility can be determined exactly, while for the next time units, there is more uncertainty (represented here by the blue bars) and the amount of flexibility certainly available (pink bars) is lower.
@@ II-G SAREF and SAREF4ENER
Smart Applications REFerence (SAREF) [^2] https://saref.etsi.org/ [SAREF] is an ontology for IoT devices and solutions published by ETSI [^3] https://www.etsi.org/ in a series of Technical Specifications initially released in 2015 and updated in 2017. Even though its initial objective was to build a reference ontology for appliances relevant to energy efficiency, SAREF can also serve as a reference model to enable better integration of data from various vertical domains in the IoT. SAREF has been extended to different domains: specifically, SAREF for Energy Flexibility (SAREF4ENER) [SAREF4ENER] is its extension for the energy domain, and aims to provide a standardized way to describe and interoperate energy-related devices, systems, and services. It allows us to describe devices such as solar panels, electric vehicles, and smart meters, as well as systems like microgrids, and energy management/demand response services. By using SAREF4ENER, developers can create applications that easily integrate and communicate with different energy devices, systems, and services, regardless of the manufacturer or protocol used, which enables more efficient and sustainable energy management.
@@
Figure 3: dCO design overview.
The World Wide Web Consortium (W3C) has developed a standard for enabling Internet of Things (IoT) interoperability, whose name is W3C Web of Things (WoT) [WoT]. In this context, Thing Descriptions (TDs) [WoT_TD] have been developed: they describe the metadata and interfaces of the considered devices that participate in the WoT. Things are organized in TD Directories (WoT TDDs), which make it possible to register, manage, and search TDs. An important feature for IoT consumers is to send queries to the WoT directory: this is achieved by having a common ontology to annotate TDs; however, before dCO, no ontologies were integrating WoT into smart buildings. dCO [DBLP:conf/esws/LaadharTP22] has been created with the purpose of achieving this integration. Specifically, dCO is an ontology designed for smart buildings, capable of representing devices and properties of the building, in addition to energy flexibility: what makes it different from other building ontologies is the integration with TDs. In this paper, we will focus on the aspect of dCO that strictly concerns energy flexibility. We now quickly describe the main modules. The core module uses the SAREF core ontology [SAREF], and enables the integration of the other modules in order to allow interoperability: it is integrated with the other modules. The TD module integrates the WoT TD ontology [TDontology] with the core module, while the devices module represents the devices. The units of measurement module use the Units of Measurement Ontology [UOMontology], while the building description module represents building-related metadata and building topology, and uses the W3C Building Topology Ontology [BOTontology]. All these modules have been described in [DBLP:conf/esws/LaadharTP22] in greater detail. Finally, there are the power, electricity and heating flow module, which is self-explanatory, and the energy flexibility module: this last is the contribution of this work, where we explain how it extends SAREF4ENER. Figure 3 shows those modules, and their main classes and properties.
@@
This section shows how SAREF4ENER has been extended to describe FOs. We show how we designed the dCO flexibility module and how we encode attributes and constraints.
@@ III-A FlexOffer Message
As mentioned in Section II, FOs are sent between actors: this is achieved by FO messages, which encode FO metadata and information about energy constraints. The attributes of an FO message describe the FO ID, state (e.g., accepted, rejected), time granularity, geographical location, time limits for acceptance and assignment, constraints on energy, time and price, and an energy schedule. Of these, the only ones to have an equivalent in SAREF4ENER are state, simple energy constraints (slice, TECs), price constraints and duration, while more advanced energy constraints (for DFOs and UFOs) and other attributes had to be added as extensions.
@@ III-B Energy Flexibility Ontological Design
Element
Mandatory
in the FO?
Exists in SAREF/
SAREF4ENER?
Corresponding class in
SAREF/SAREF4ENER
Notes
Standard FlexOffer
ID Yes No
State Yes Yes PowerSequenceState
StateReason No No
NumSecondsPerInterval Yes No
CreationTime Yes No
CreationInterval Yes No
OfferedByID Yes No
LocationID No No In SAREF, a device can have a location
AcceptBeforeTime No No
AcceptBeforeInterval No No
AssignmentBeforeTime No No
AssignmentBeforeInterval No No
StartAfterInterval No No
StartBeforeTime Yes No
StartBeforInterval No No
EndAfterTime No No
EndAfterInterval No No
EndBeforeTime No No
EndBeforeInterval No No
StartAfterTime Yes No
FlexOfferProfileConstraints Yes No
FlexOfferPriceConstraint No No
DefaultSchedule No No
FlexOfferSchedule No No
FlexOfferProfileConstraints
EnergyConstraintsList Yes Yes Energy
SAREF includes EnergyMax
and EnergyMin
PriceConstraint No Partially hasPrice
SAREF does not include min
and max price
MinDuration No Yes ActiveDurationMin
MaxDuration No Yes ActiveDurationMax
TotalCostConstraint No No
ScheduleSlice attributes:
Duration No Yes DefaultDuration
EnergyAmount Yes Yes EnergyExpected
Price No No
TABLE I: FlexOffer mapping to SAREF4ENER - core message.
Element
Mandatory
in the FO?
Exists in SAREF/
SAREF4ENER?
corresponding class in
SAREF/SAREF4ENER
Notes
Total energy FlexOffer
TotalEnergyConstraint
(lower and upper)
No Yes Energy
SAREF includes EnergyMax
and EnergyMin
Dependency FlexOffer
DependencyEnergyConstraintList No No
PriceConstraint
(minPrice, maxPrice)
No Partially hasPrice
MinDuration No Yes ActiveDurationMin
MaxDuration No Yes ActiveDurationMax
Uncertain FlexOffer
UncertainFunctions No No
UncertainThreshold No No
MinDuration No Yes ActiveDurationMin
MaxDuration No Yes ActiveDurationMax
TABLE II: FlexOffer mapping to SAREF4ENER.